# On the Geometry of Semantics in Next-token Prediction

### v1

@@ F.1 Semantics for Individual Concept
In this section we include the plots illustrating that individual concept do not contain interpretable semantics as discussed in sec 5
@@
F.2 Experiment on the Semantics of Pretrained Context Embeddings
To identify the semantics in context embeddings from pretrained transformers, we chose GPT-2 Radford et al. [2019] and extracted $50,000$ text sequences, each 11 tokens in length from GPT-2’s training corpus, the OpenWebText dataset. By segmenting these sequences into contexts of varying lengths (1-10 tokens), we generated $500,000$ contexts, which we designate as the "simplified OpenWebText" dataset.
To obtain concepts from pretrained models, instead of performing SVD on $\widetilde{\bm{S}}$ (which is computationally expensive to construct and factorize), we work directly with the embeddings of the pretrained model. For word concepts, we extract ${\bm{W}}$ from GPT-2’s decoder and take the matrix $\bm{U}$ of word concepts to be its left singular matrix. For context concepts, we form a $d\times m$ matrix ${\bm{H}}$ by concatenating the embeddings (last-MLP layer representations before the decoder) of GPT-2 for each context in our subset. This matrix corresponds to only a portion of the “true” embedding matrix for the entire dataset. To ensure the coupling between ${\bm{W}}$ and ${\bm{H}}$ that is inherent to the SVD of $\widetilde{\bm{S}}$ (which we don’t have direct access to), we compute the matrix ${\bm{V}}$ of context concepts as ${\bm{V}}={\bm{H}}^{\top}\mathbf{R}\bm{\Sigma}^{-1/2}$, where $\bm{U}\bm{\Sigma}^{1/2}\mathbf{R}$ denotes the SVD of ${\bm{W}}$ that we already computed.
We include the orthant-clustering result on GPT2’s pretrained word embedding with $p=4,5,6$. (Fig. 15, 16, 17)
Figure 15: orthant-based clustering on Pretrained word embedding, $p=4$, GPT2
Figure 16: Semantics on Pretrained word embedding, $p=5$, GPT2
Figure 17: orthant-based clustering on Pretrained word embedding, $p=6$, GPT2
F.3 Rate of learning
@@
F.4 $d>V$
Fig. 19 shows the evolution of singular values for squared loss when the network’s embedding dimension is smaller than the vocabulary size, i.e., $d<V$, as described in Section A.1.
Figure 19: evolution of the learned logit matrix’s singular values during training for $3=d<V$, note that the 3 singular values are converging to the $d$ largest singular values
F.5 Semantics from K-means Spectral Clustering on Simplified TinyStories and Wiki-Text
Please see Fig. 20 and 21.
Figure 20: 32-means spectral clustering, $p=5$, Simplified TinyStories
Figure 21: 64-means spectral clustering, $p=6$, Simplified WikiText
F.6 More orthant-based clustering result on Simplified TinyStories and Wiki-Text
We include some examples of orthant-based clustering results in Fig. 22 to 25. We selected $p=4$, $p=5$ for Simplified TinyStories and $p=5$, $p=6$ for Simplified WikiText because we observe the most human-interpretable concepts.
Figure 22: Orthant-based clustering, $p=4$, Simplified TinyStories
Figure 23: Orthant-based clustering, $p=5$, Simplified TinyStories
Figure 24: Orthant-based clustering, $p=5$, Simplified Wiki-Text
Figure 25: Orthant-based clustering, $p=6$, Simplified Wiki-Text

### v2

@@ D.4 Semantics for Individual Concept
We include the plots illustrating that individual concept do not contain interpretable semantics as discussed in Sec 4.
@@
D.5 Rate of learning
@@
D.6 Extension to Masked Language Models
To explore semantic structure in model trained with different objectives, we applied orthant-based clustering to token embeddings from a pretrained BERT model. As shown in Fig. 14, the method recovers a range of interpretable semantic categories.
3-digit Number
Historical Years
Action
Administration
Given Names
Operations
Name Suffixes
Verb Suffixes
Figure 14: Semantics identified by orthant-based clustering on BERT’s word embeddings. In BERT’s tokenizer, tokens prefixed with ## represent subword units that appear within or at the end of a word.
D.7 Extension to Higher-Dimensional Embeddings
As discussed in Sec. 4.2, we extend our analysis to the Qwen-7B model. We summarize representative semantic categories for selected orthant configurations in Table 1.
Table 1: Examples of semantic categories recovered from Qwen-7B using orthant-based clustering on its 4096-dimensional word embeddings. Each row corresponds to a specific sign configuration $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ along the top singular concept directions.
D.8 Semantics from K-means Spectral Clustering on Simplified TinyStories and Wiki-Text
Unlike orthant-based clustering, which can produce up to $2^{p}$ clusters when selecting $p$ top concepts, vanilla k-means on principal components results in $p$ clusters. Thus, when evaluating spectral methods, we choose $k$ on the order of $2^{p}$ for fair comparison. Specifically We apply $k$-means clustering using pairwise distances of analyzer vectors restricted to their top $p=\log_{2}(k)$ dimensions. As mentioned, this choice corresponds to the top-$p$ concept directions anddiffers from vanilla spectral clustering, which selects $k$ principal components. We chose $k=32$ for Simplified TinyStories and $k=64$ for WikiText-2. Fig. 15 displays example clusters.
Past-tense verbs Refer to caption
Proper names Refer to caption
Numbers Refer to caption
Conjunctions Refer to caption
Appendix E Use of AI Assistants
We used AI assistants (ChatGPT, Claude, and GitHub Copilot) to support manuscript preparation. ChatGPT and Claude assisted with text editing, grammar correction, and phrasing refinement. GitHub Copilot assisted with code generation and debugging. All conceptual contributions, theoretical results, experimental designs, and scientific conclusions are solely the work of the authors. AI-generated content was critically reviewed by the authors.